Grade 11Physical Chemistry

the empirical formula of a compound formed when 7. 30 grams of iron powder reacts completely with 6. 30 grams of powdered sulphur is? please show full method

ashu
12 Years agoGrade 11

1 Answer

Sunil Kumar FP
12 Years ago
mole of iron =7.3/56=.13
mole of sulfur formed=6.3/32=.196
ratio simplest=2:3
therefore empirical formula =fe2s3

thanks and regards
sunil kumar
ask IItian faculty